# On the second Bohr radius for vector valued pluriharmonic functions

### v1

@@ 3.2. Second Bohr radii for homogeneous plynomials
$\sum_{|\alpha|=m}\left\lVert\left(\frac{a_{\alpha}-b_{\alpha}}{2}\right)\right\rVert^{p}\leq 2^{p}\,(\pi_{1}(I_{X}))^{2-p}\,\left(t\,n^{\frac{1}{q}}\right)^{(2-p)m}(\chi_{M}(\mathcal{P}(^{m}Z)))^{2-p}\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{2(p-1)m}.$
@@
$\left(\sum_{|\alpha|=m}\left\lVert a_{\alpha}\right\rVert^{p}\right)^{\frac{1}{p}}\leq 4\,(\pi_{1}(I_{X}))^{\frac{(2-p)}{p}}\,\left(t\,n^{\frac{1}{q}}\right)^{\frac{(2-p)m}{p}}(\chi_{M}(\mathcal{P}(^{m}Z,Y)))^{\frac{(2-p)}{p}}\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{\frac{2(p-1)m}{p}}.$
@@
$\leq 2^{2p+1}(\pi_{1}(I_{X}))^{2-p}\,r^{pm}\,s^{pm}\,\left(t\,n^{\frac{1}{q}}\right)^{(2-p)m}(\chi_{M}(\mathcal{P}(^{m}Z,Y)))^{2-p}\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{2(p-1)m},$
@@
$L^{m}_{\lambda}(B_{Z},p,X)\geq E_{2}(X)\,\,\frac{\lambda^{\frac{1}{m}}}{s\,\left(t\,n^{\frac{1}{q}}\right)^{\frac{(2-p)}{p}}(\chi_{M}(\mathcal{P}(^{m}Z)))^{\frac{(2-p)}{pm}}\left\lVert I:\ell^{n}_{2}\rightarrow Z\right\rVert^{\frac{2(p-1)}{p}}},$
@@
$L^{m}_{\lambda}(B_{Z},p,X)\geq\frac{E_{2}(X)\,\lambda^{\frac{1}{m}}}{s\,\left(t\,n^{\frac{1}{q}}\right)^{\frac{(2-p)}{p}}}\max\left\{\frac{1}{(\sqrt{n})^{1-\theta}\,\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert},\,\frac{\left\lVert I:\ell^{n}_{2}\rightarrow Z\right\rVert^{-\theta}}{\left(\frac{m^{m}}{m!}\right)^{(1-\theta)/m}\,\left\lVert Id:Z\rightarrow\ell^{n}_{1}\right\rVert^{1-\theta}}\right\},$

### v2

@@ 3.2. Second Bohr radii for homogeneous plynomials
$\sum_{|\alpha|=m}\left\lVert\left(\frac{a_{\alpha}-b_{\alpha}}{2}\right)\right\rVert^{p}\leq 2^{p}\,(\pi_{1}(I_{X}))^{2-p}\,\left(t\,n^{\frac{1}{q}}\right)^{(2-p)m}(\chi_{M}(\mathcal{P}(^{m}Z)))^{2-p}\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{2(p-1)m}.$
@@
$\left(\sum_{|\alpha|=m}\left\lVert a_{\alpha}\right\rVert^{p}\right)^{\frac{1}{p}}\leq 4\,(\pi_{1}(I_{X}))^{\frac{(2-p)}{p}}\,\left(t\,n^{\frac{1}{q}}\right)^{\frac{(2-p)m}{p}}(\chi_{M}(\mathcal{P}(^{m}Z,Y)))^{\frac{(2-p)}{p}}\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{\frac{2(p-1)m}{p}}.$
@@
$\leq 2^{2p+1}(\pi_{1}(I_{X}))^{2-p}\,r^{pm}\,s^{pm}\,\left(t\,n^{\frac{1}{q}}\right)^{(2-p)m}(\chi_{M}(\mathcal{P}(^{m}Z,Y)))^{2-p}\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{2(p-1)m},$
@@
$L^{m}_{\lambda}(B_{Z},p,X)\geq E_{2}(X)\,\,\frac{\lambda^{\frac{1}{m}}}{s\,\left(t\,n^{\frac{1}{q}}\right)^{\frac{(2-p)}{p}}(\chi_{M}(\mathcal{P}(^{m}Z)))^{\frac{(2-p)}{pm}}\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{\frac{2(p-1)}{p}}},$
@@
$L^{m}_{\lambda}(B_{Z},p,X)\geq\frac{E_{2}(X)\,\lambda^{\frac{1}{m}}}{s\,\left(t\,n^{\frac{1}{q}}\right)^{\frac{(2-p)}{p}}}\max\left\{\frac{1}{(\sqrt{n})^{1-\theta}\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert},\,\frac{\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{-\theta}}{\left(\frac{m^{m}}{m!}\right)^{(1-\theta)/m}\,\left\lVert\mathrm{Id}:Z\rightarrow\ell^{n}_{1}\right\rVert^{1-\theta}}\right\},$